OpenStudy (anonymous):

The x-intercept of the graph of f(x)=3log(x+5)+2 is: a. 10^-2/3 - 5 b. 10^-2/3 + 5 c. 10^2/3 + 5 d. 10^2/3 - 5

OpenStudy (anonymous):

for x - intercept put f(x) = 0 and solve for x

OpenStudy (anonymous):

I got -5.6 for x...

OpenStudy (anonymous):

it should be (A)

OpenStudy (anonymous):

-4.78

OpenStudy (anonymous):

when y=0 3log(x+5)+2=0 log(x+5)= -2/3 10^log(x+5) = 10^-2/3 therefore, (x+5) = 10^-2/3 thus, x=10^-2/3 - 5

OpenStudy (anonymous):

did you understand how its done?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes thank you